OpenStudy (anonymous):

y = -2x + 19 y = x + 7

OpenStudy (jojo4eva):

y=x+19y-7

OpenStudy (anonymous):

x+7=-2x+19 3x=12 x=4

OpenStudy (jojo4eva):

yes thank you @Principes i didnt want them to have the wrong answer

OpenStudy (anonymous):

there is not a answer its impossible.